OpenStudy (anonymous):

When finding the sum of the arithmetic sequence 4 + 1 + -2 + -5 + -8 + -11, you will have a step of work that contains the number -3. Use complete sentences to describe how the -3 was calculated and what it means.

OpenStudy (anonymous):

I don't understand this could someone walk me through this?

OpenStudy (amistre64):

what are the steps they say youhave to take to find the sum of the sequence?

OpenStudy (amistre64):

my approach may not be their approach so my step with a -3 in it may not be their step with a -3 in it, or i may have many steps with a -3 in it which do i choose?

OpenStudy (anonymous):

I am confused?

OpenStudy (amistre64):

an arith sequence is generated by adding a common number to each term to get the next term. if we start with a first term, ao and add a constant value, d, to it each time w generate the sequence: ao = ao = ao + 0d a1 = ao + d = ao + 1d a2 = a1 + d = ao + 2d a3 = a2 + d = ao + 3d .... as a result an = ao + nd when we add all these up, we have n+1 terms, 0 to n ao + 0d ao + 1d ao + 2d ao + 3d ... ao + nd ---------- A + B where A = ao(n+1) and B=d(0+1+2+3+...+n)

OpenStudy (amistre64):

lost connect .. reconnecting ... laglaglag ....

OpenStudy (anonymous):

Okay so looking at the work you provided @amistre64 I don't see a -3. @iGreen @Callisto @phi do any of you think you can help me by finishing what amistre started?

OpenStudy (phi):

4 + 1 + -2 + -5 + -8 + -11 if you were going to add up those numbers, what would you do ? What would be your first step?